BD L 60E VA - pino O roller

9. [1pt]
Two forces are applied to the truss as shown, F1 = 182 N and F2 = 187 N. The truss is supported by a pin joint at A and a roller at E. Assume L = 2.60 m.
Indicate whether the member AB is in compression (C) or in tension (T). You only have 1 try!

Correct, computer gets: T

10. [1pt]
Calculate the value of the force in member CD. Enter a negative value if the member is in compression.

Correct, computer gets: -374.0 N

11. [1pt]
Calculate the value of the force in member AD. Enter a negative value if the member is in compression.

Answer:  Last Answer: 416.99 N
Incorrect, ONE try left!!

12. [1pt]
Calculate the value of the force in member AB. Enter a negative value if the member is in compression.

Answer:  Last Answer: 83.13 N

**Please solve 11 and 12
Answer #1

Please give ratings

187 N L= 2.60m 30 (90-60) = 30 - 182N B 60° 60E RAH TRA * Pin- Supports provide vertical of well as horizontal reaction * Rol

E Ex=0 >RAH = 369N = RAH = 182+187 Fy 20 => RAN=REV Now considering moment about A. EMATO 182 x 2.6 +187* 5.2 = REV (3.002) 4